# An explicit construction of Kaleidocycles by elliptic theta functions

### v3

@@ Abstract.
We consider the configuration space of ordered points on the two-dimensional sphere that satisfy a specific system of quadratic equations. We construct periodic orbits in this configuration space using elliptic theta functions and show that they simultaneously satisfy semi-discrete analogues of mKdV and sine-Gordon equations. The configuration space we investigate corresponds to the state space of a linkage mechanism known as the Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle. A key consequence of our construction is the proof that Kaleidocycles exist for any number of tetrahedra greater than five. Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system is explicitly solved to generate an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
@@
The first-named author was partially supported by KAKENHI, Grant-in-Aid for Scientific Research (B) 25K00921 and (S) 25H00399. The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229.
@@ 1. Introduction
Given a natural number $k$ and a real number $1<\cos\lambda<1$, we consider the following space of configurations of $k+1$ ordered points on the 2-sphere $S^{2}$ satisfying certain quadratic constraints:
@@
Our main result states that for any $k\geq 6$, there exists $\lambda$ such that a circle can be embedded in $\mathcal{C}^{\pm}_{k,\lambda}$. Moreover, we explicitly construct a circle in terms of elliptic theta functions. This result implies the existence of a particular class of linkage mechanisms called Kaleidocycles (Fig. 1 (Left)). These mechanisms consist of $k$ identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints, forming a ring. Although Kaleidocycles have been studied by several authors [5, 9, 30, 31], the rigorous proof of their existence beyond trivial ones has remained an open problem. Since an element of $\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a Kaleidocycle with $k$-tetrahedra in a specific position (Section 2.1) and its motion is described by a path in the configuration space, our main result provides an explicit construction and its periodic motion for Kaleidocycles of any $k\geq 6$.
@@
The key observation is to identify a Kaleidocycle with a closed spatial polygonal curve formed by connecting the midpoints of the hinge edges. Then, the angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the signed curvature angle (§2.1). An element $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a sequence of binormals. In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle $\lambda$. In our construction, it plays the central role to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
@@
It was known in [21] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations are defined on the configuration space (see §3.2). The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [12]5). Moreover, the sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 18].
@@
Both views combined, a flow generated by an integrable system on a real-algebraic set $\mathcal{C}^{\pm}_{k,\lambda}$ is explicitly integrated using elliptic theta functions. Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, the Kaleidocycle.
@@
The structure of the paper is as follows. Section 2 introduces the definition of Kaleidocycles. They are identified with a certain class of discrete spatial curves. As essential invariants to describe the state of a Kaleidocycle, the signed curvature angle and torsion angle are defined together with the Frenet frame. Section 3 reviews the connection between the theory of curves and integrable systems. The construction of discrete spatial curves and their Frenet frames in terms of the $\tau$ functions is presented. A previous result that describes the motion of a Kaleidocycle using semi-discrete integrable equations is recalled. Section 4 constructs a set of $\tau$ functions with three parameters $v,r,y$ using elliptic theta functions. The resulting curve is shown to have a constant torsion angle. In Section 5, a family of time-parametrised curves is constructed by introducing an additional parameter $t$. We then study the time evolution equations satisfied by the vertex positions and curvature angle of the curve and show connections to previous research. Section 6 investigates a necessary and sufficient condition for the resulting curve to be closed for every $t$. We prove there exists a set of parameters that satisfies the closure condition for a given $k\geq 6$, which proves the existence of Kaleidocycles. Section 7 provides examples of Kaleidocycles and corresponding semi-discrete surfaces with figures. We present a conjecture (Conjecture 7.1) suggested by numerical experiments: the Kaleidocycles constructed in this paper seem to be so-called Möbius Kaleidocycles.
@@ 2.1. Configuration space of a Kaleidocycle
A Kaleidocycle consists of (typically six) identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints to form a ring. It can be folded from a sheet of paper and is an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid. The origin of Kaleidocycle is obscure, but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R. For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [30] and the references therein). However, there is no literature that discusses the existence of such objects for arbitrary $k$.
@@
The solution space of (2.1) admits an action of the orthogonal group $O(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $3\cdot 2k-5k-3-\dim(O(3))=k-6$, where $5k$ is the number of edges and $3$ comes from the equation $x_{0}+x_{1}=0$. This simple counting estimate is often called Maxwell’s degree-of-freedom counting law [9]. However, this estimate is sometimes violated in interesting examples. The Bricard 6R is a case of $k=6$. Its configuration space has dimension one, which is greater than the expected dimension of 0, giving an example of an overconstrained system  [9]. When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [22]. [^1] Some of the findings of [22] are also published in [31]. For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of an underconstrained linkage mechanism whose configuration space has a smaller dimension than the expected one.
@@
Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k)$ be the mid-point of the hinge edge. There exists $l>0$ such that $l|b_{n+1}\times b_{n}|=|\gamma_{n+1}-\gamma_{n}|$. By 2.2, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $b_{n}$ and $b_{n+1}$. Therefore, $b_{n+1}\times b_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, and we have $l\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$. This means $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$.
Conversely, given $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$, define $\gamma_{n}=l\sum_{m=0}^{n-1}(b_{m+1}\times b_{m})$, where $l$ is the distance of the opposing hinge edges of the tetrahedron. We set $x_{2n}=\gamma_{n}-\frac{l_{01}}{2}b_{n}$ and $x_{2n+1}=\gamma_{n}+\frac{l_{01}}{2}b_{n}$. Then, $(x_{0},\ldots,x_{2k+1})$ satisfies Equation 2.1.
It is clear that the above constructions are mutually inverse. ∎
@@ 2.2. Discrete spatial curve of a constant torsion angle
As in the case of a smooth curve, a discrete spatial curve of a unit segment length is determined by its torsion and curvature angles. Therefore, the motion of a discrete curve of a constant torsion angle is determined by the set of functions $\kappa_{n}(t)$, which can be freely chosen.
In this language, a Kaleidocycle is equivalent to a discrete closed ($\gamma_{k}=\gamma_{0})$ spatial curve with constant torsion angle $\lambda$ determined by the binormal $\widetilde{B}_{n}=b_{n}$. Our approach is to utilise this identification.
The fact that a Kaleidocycle is closed enforces interesting topological constraints. For example, the self-linking number must be a half-integer, and hence, stay constant during the motion. By the Călugăreanu-Fuller-White theorem [1], the writhing number of the curve must remain constant as well. The self-linking number also tells that the configuration space of a Kaleidocycle is not generally connected.
@@ 3. Curves and integrable systems
There are deep connections between differential geometry and the theory of integrable systems, where the integrable differential or difference equations arise as compatibility conditions of the geometric objects (see, for example, [29]). In the context of the deformation of spatial and plane curves, the Frenet frame and its deformation are described by the system of linear partial differential equations. The compatibility condition yields the so-called Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example. In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [10, 25, 29]. The connections are also investigated in a discrete setting, where curves are discretised as piecewise linear (polygonal) curves. For example, continuous deformations of discrete plane and spatial curves have been considered in [6], where the deformations are governed by the semi-discrete analogue of the mKdV and the NLS equations. Discrete deformations of the discrete plane and spatial curves have been studied in [18]. In principle, for a given solution to the nonlinear differential/difference equation for the curvature and torsion, one can reconstruct the curve by solving the associated system of linear differential/difference equations for the Frenet frame and integrating the tangent vector.
@@ 3.2. Motion of a Kaleidocycle on the osculating plane
(3.24) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\tilde{\rho}^{2},\quad|{\gamma}_{n+1}-{\gamma}_{n}|=1\quad(0\leq\forall n\leq k),$
@@
Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle in Theorem 4.2.
@@ 4.1. A family of curves with a constant torsion angle
This theorem follows as a corollary of Theorem 5.1 and will be proven in §5.1.
@@
Note that $2\tan\frac{\kappa_{n}}{2}$ is often considered as the discritised curvature (see, for example, [14]), which is expressed using Jacobi $\mathrm{cn}$-function as follows:
@@ 5. Deformation of a discrete curve with a constant torsion angle
By introducing the deformation parameter $t$ to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem $5.1$). We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature (§5.2, §5.3, §5.4). Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [18] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [21] are discussed. This is also summarised in Table 1 in §6.
5.1. Explicit formula of the curve with deformation parameter
@@
(5.3) $\begin{split}\kappa_{n}\left(t\right)&=\kappa_{n}\left(t;v,y\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta_{1}F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right).\end{split}$
@@
From this we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows. Theorem 4.2 follows from Theorem 5.1 by setting $t=0$. ∎

### v4

@@ Abstract.
We study a configuration space consisting of ordered points on the two-dimensional sphere satisfying a system of quadratic constraints. We construct explicit periodic orbits in the configuration space using elliptic theta functions. The constructed orbits simultaneously satisfy semi-discrete analogues of the modified KdV and sine-Gordon equations. This configuration space arises as the state space of a linkage mechanism called a Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle. A key consequence of our construction is the proof that Kaleidocycles exist for any number of tetrahedra $k\geq 6$. Our approach is founded on the relationship between the deformation of spatial curves and integrable systems. The motion of the mechanism is interpreted as a deformation of a closed discrete spatial curve with constant torsion angle. This provides an explicit example in which an integrable system is solved to generate periodic orbits in a real solution set of polynomial equations arising from geometric constraints.
@@
The first-named author was partially supported by JSPS KAKENHI Grant Numbers JP25K00921 and JP25H00399. The second-named author was partially supported by JSPS KAKENHI Grant Number JP25K21661. The third-named author was partially supported by JSPS KAKENHI Grant Numbers JP25K00921 and JP25K17297.
@@ 1. Introduction
Given a natural number $k$ and a real number $0<\lambda<\pi$, we consider the following configuration space of $k+1$ ordered points on the 2-sphere $S^{2}$ satisfying certain quadratic constraints:
@@
This configuration space naturally appears as the state space of a linkage mechanism called a Kaleidocycle. A Kaleidocycle consists of $k$ identical equifacial tetrahedra connected by hinge joints along pairs of opposite edges so that they form a ring (Fig. 1 (Left)). The mechanism exhibits a characteristic turning motion. Although Kaleidocycles have been studied by several authors [5, 9, 30, 31], the rigorous proof of their existence beyond trivial cases has remained an open problem.
Our main result (Theorem 6.1) states that for any $k\geq 6$ there exists $\lambda$ for which the configuration space $\mathcal{C}^{\pm}_{k,\lambda}$ contains an embedded circle. We explicitly construct this circle using elliptic theta functions. Since a point of $\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a particular configuration of the mechanism, a closed path in this space describes a periodic motion of the Kaleidocycle. Therefore our result provides a constructive proof of the existence of Kaleidocycles consisting of $k$ tetrahedra for all $k\geq 6$.
@@
The key observation is that a Kaleidocycle can be identified with a closed discrete spatial curve obtained by connecting the midpoints of the hinge edges. In this interpretation the hinge edges correspond to the binormal directions of the curve. The angles between consecutive hinge edges correspond to a torsion angle, while the turning angles of the polygon correspond to signed curvature angles (Section 2.1). Hence an element $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$ can be regarded as a sequence of binormal vectors of a closed polygonal curve with constant torsion angle $\lambda$. In this geometric formulation the motion of a Kaleidocycle is modelled as a deformation of a closed polygonal curve preserving both the segment length and the torsion angle. In our construction, the compatibility conditions of the discrete Frenet frame describing such a deformation are compared with the functional identities of elliptic theta functions.
@@
It was known in [20] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations are defined on the configuration space (Section 3.2). The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [12] (Section 5). Moreover, the sweeping surface of the motion corresponds to a semi-discrete analogue of the fully-discrete constant negative curvature surface (K-surface) studied in [2, 18].
@@
Combining both views, we explicitly integrate the flow generated by an integrable system on a real-algebraic set $\mathcal{C}^{\pm}_{k,\lambda}$ using elliptic theta functions. Our work connects seemingly distant fields of integrable systems and the kinematics of linkages via an interesting object, the Kaleidocycle.
@@
The structure of the paper is as follows. Section 2 introduces Kaleidocycles and their description as discrete spatial curves. The signed curvature angle and torsion angle are defined together with the associated Frenet frame. Section 3 reviews the relationship between discrete curves and integrable systems through the $\tau$ functions, and recalls previous results describing Kaleidocycle motion by semi-discrete equations. Section 4 constructs a set of $\tau$ functions that lead to discrete curves with constant torsion angle using elliptic theta functions. Section 5 introduces a time parameter and derives the evolution equations satisfied by the curve. Section 6 proves our main result (Theorem 6.1) that for every $k\geq 6$ there exist parameters for the $\tau$ functions for which the curve closes, establishing the existence of Kaleidocycles. Finally, Section 7 presents numerical examples and a conjecture that the constructed solutions correspond to the so-called Möbius Kaleidocycles.
We briefly explain the strategy of the proof of the main result, Theorem 6.1, which establishes the existence of Kaleidocycles for every $k\geq 6$. The argument consists of three main steps. First, in 2.5, we identify Kaleidocycles with closed discrete curves of constant torsion angle. Next, in Theorem 5.1, we construct a family of motions of curves having the desired geometric properties. Finally, in Section 6, we determine a choice of parameters for which these curves become closed.
@@ 2.1. Configuration space of a Kaleidocycle
A Kaleidocycle consists of (typically six) identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints to form a ring. It can be folded from a sheet of paper and is an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid. The origin of Kaleidocycles is obscure, but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R. For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [30] and the references therein). However, there is no literature that discusses the existence of such objects for arbitrary $k$.
@@
The solution space of (2.1) admits a natural action of the orthogonal group $O(3)$ of $\mathbb{R}^{3}$. A naive dimension count then predicts that the quotient space should have dimension
$3\cdot 2k-5k-3-\dim(O(3))=k-6,$
where $5k$ is the number of edges and the term $3$ comes from the constraint $x_{0}+x_{1}=0$. This heuristic is commonly referred to as Maxwell’s degree-of-freedom counting law [9].
It is well known, however, that this estimate can fail in nontrivial ways. A classical example is the Bricard 6R linkage ($k=6$), which is equivalent to the classical $6$-Kaleidocycle. Its configuration space is one-dimensional, exceeding the expected dimension 0, and hence it provides a prototypical example of an overconstrained mechanism [9].
For $k\geq 7$, numerical evidence indicates that, when the angle between adjacent hinge edges attains a critical value, the configuration space degenerates to a circle. The existence of this single degree of freedom was first identified by the first-named author together with collaborators, leading to a subsequent patent application [19]. Shortly thereafter, related results were disseminated in the literature [31].
For $k>7$, this family of Kaleidocycles–referred to as Möbius Kaleidocycles [19]–thus provides a rare example of an underconstrained linkage, whose configuration space has strictly smaller dimension than that predicted by Maxwell’s count.
@@
Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k)$ be the mid-point of the hinge edge. By 2.2, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $b_{n}$ and $b_{n+1}$. Therefore, $b_{n+1}\times b_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, and we have $l(b_{n+1}\times b_{n})=\gamma_{n+1}-\gamma_{n}$ for some $l\in\mathbb{R}$. Then, $l\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$. This means $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$.
Conversely, given $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$, define $\gamma_{n}=\frac{l}{\sin\lambda}\sum_{m=0}^{n-1}(b_{m+1}\times b_{m})$, where $l$ is the distance of the opposing hinge edges of the tetrahedron. We set $x_{2n}=\gamma_{n}-\frac{l_{01}}{2}b_{n}$ and $x_{2n+1}=\gamma_{n}+\frac{l_{01}}{2}b_{n}$. Then, $(x_{0},\ldots,x_{2k+1})$ satisfies Equation 2.1.
It is clear that the above constructions are mutually inverse and $O(3)$ equivariant. ∎
@@ 2.2. Discrete spatial curve of a constant torsion angle
As an immediate corollary of 2.3, we obtain the following identification.
Proposition 2.5.
A state of a Kaleidocycle is identified with a closed discrete spatial curve $(\gamma_{k}=\gamma_{0})$ of constant torsion angle $\lambda$ whose binormal vectors $\widetilde{B}_{n}$ satisfy (1.1) with $b_{n}=\widetilde{B}_{n}$.
As in the smooth case, a discrete spatial curve of unit segment length is determined by its curvature and torsion angles. Accordingly, the motion of a discrete curve with constant torsion angle is governed by the family of curvature functions $\kappa_{n}(t)$.
The fact that a Kaleidocycle is closed imposes interesting topological constraints. The signed curvature angles $\kappa_{n}(t)$ are therefore subject to non-trivial restrictions. In particular, the self-linking number is a half-integer, and hence remains constant during the motion. By the Călugăreanu–Fuller–White theorem [1], the writhe of the curve must also remain constant. Moreover, the self-linking number shows that the configuration space of a Kaleidocycle is not connected in general.
@@ 3. Curves and integrable systems
There are deep connections between differential geometry and the theory of integrable systems, where the integrable differential or difference equations arise as compatibility conditions of the geometric objects (see, for example, [29]). In the context of the deformation of spatial and plane curves, the Frenet frame and its deformation are described by the system of linear partial differential equations. The compatibility condition yields the so-called Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example. In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [10, 25, 29]. The connections are also investigated in a discrete setting, where curves are discretized as piecewise linear (polygonal) curves. For example, continuous deformations of discrete plane and spatial curves have been considered in [6], where the deformations are governed by the semi-discrete analogue of the mKdV and the NLS equations. Discrete deformations of the discrete plane and spatial curves have been studied in [18]. In principle, for a given solution to the nonlinear differential/difference equation for the curvature and torsion, one can reconstruct the curve by solving the associated system of linear differential/difference equations for the Frenet frame and integrating the tangent vector.
@@ 3.2. Motion of a Kaleidocycle on the osculating plane
(3.24) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\rho^{2},\quad|{\gamma}_{n+1}-{\gamma}_{n}|=1\quad(0\leq\forall n\leq k),$
@@
Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle in 4.2.
@@ 4.1. A family of curves with a constant torsion angle
This proposition follows as a corollary of Theorem 5.1 and will be proven in §5.1. Note that $\widetilde{G}$ and $G$ have the same simple zeros. Therefore, (4.12) remains valid even at points where $\widetilde{G}$ vanishes. Moreover, it follows from the zero distribution of the theta functions that both $\beta_{1}$ and $F_{n}$ are non-vanishing.
Combining this with (3.1), it follows that (3.10) and (4.13) hold for any $v$,$r$,$y$, and $n$.
@@
Note that $2\tan\frac{\kappa_{n}}{2}$ is often considered as the discretized curvature (see, for example, [15]), which is expressed using Jacobi $\mathrm{cn}$-function as follows:
@@ 5. Deformation of a discrete curve with a constant torsion angle
By introducing the deformation parameter $t$ to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem $5.1$). We then derive the evolution equation for the vertex positions of the curve in Theorem 5.85.2), which reduces to the previously known equations presented in §3.2 under the specific specialization of the global motion of the curve, as detailed in Theorem 5.11. Furthermore, we derive the evolution equations for the (potential of the) curvature in Theorem 5.16 and Theorem 5.195.3, §5.4). Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [18] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [20] are discussed. This is also summarized in Table 1 in §6.
5.1. Explicit formula for the curve with deformation parameter
@@
(5.3) $\begin{split}\kappa_{n}\left(t\right)&=\kappa_{n}\left(t;v,y\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta_{1}F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right),\end{split}$
and has a period in $t$:
(5.4) $\kappa_{n}\left(t+y\right)=-\kappa_{n}\left(t\right).$
@@
From this we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows. Periodicity of the signed curvature angle (5.4) follows from (A.12). 4.2 follows from Theorem 5.1 by setting $t=0$. ∎